# A Mind Cannot Be Smeared Across Time

### v1

@@ Abstract
Whether machines can be conscious depends not only on what they compute, but when they compute it. Most deployed artificial systems realise their functions via sequential or time-multiplexed updates. Conscious experience appears unified and simultaneous. I show that this difference matters formally. I augment Stack Theory with algebraic laws relating within time-window constraint satisfaction to conjunction. I introduce a precise temporal semantics over windowed trajectories $\tau^{\Delta,s}$ and prove that existential temporal realisation $\Diamond_{\Delta}$ does not preserve conjunction. A system can realise all the ingredients of experience across time without ever instantiating the experienced conjunction itself. I then distinguish two postulates. StrongSync requires objective co-instantiation of the grounded conjunction within the window, while WeakSync permits temporal “smearing”. I formalise concurrency-capacity to measure what is needed to satisfy StrongSync. Finally, I review neurophysiological evidence suggesting that consciousness depends on phase synchrony and effective connectivity, and that loss of consciousness is often associated with its breakdown. This evidence makes WeakSync less plausible. Under StrongSync, software consciousness on strictly sequential substrates is impossible for contents whose grounding requires two or more simultaneous contributors. The more parts from which simultaneous contribution required, the more concurrency capacity is required. The hardware matters. Consciousness attribution therefore requires architectural inspection, not just functional performance.
@@ 1 Introduction
A central open question in machine consciousness is whether behaviourally and functionally matched systems must be matched in phenomenal realisation. Stack Theory seeks to approach this question by formalising cognitive processes in terms of abstraction layers [undeff, undefb, undefh]. Various aspects of incumbent theories of consciousness are formalised [undefw, undefj, undefn, undefv, undefy, undef], organised into a hierarchy of causal identities [undefb], and derived as a necessary consequence of a generalisation-optimal learning theory that makes valence primary [undefe, undefg, undefc, undefa]. Those necessary ingredients of consciousness are beyond the scope of this paper. Here I am concerned with a question of sufficiency. Crucially, Stack Theory shows how the behaviour of two systems can appear identical at one level of abstraction, while being different in the levels below. This means that conscious experience can be purely functional, but still depend on more than surface level function. It can depend not just on what is computed, but when and how. The ingredients of what looks like a single experienced moment at a high level of abstraction may be smeared across objective micro-time at a lower level of abstraction, such that no objective time-slice contains the full grounded conjunction (meaning the parts of a subjective moment of consciousness exist at different objective times, never together). If the ingredients of a subjective moment need not synchronise in objective time for that moment to be experienced, then current hardware is sufficient to host machine consciousness. That also means a “liquid-brained” swarm intelligence [undefx] like an ant colony or a population of humans can be conscious. However, if synchronised broadcast and causal co-dependence matter, then far fewer systems can be conscious. The hardware matters, because the parts of a subjective conscious moment need to be connected in objective time to be experienced. This question of architectural sufficiency is called The Temporal Gap.
The Temporal Gap was proposed in earlier work [undeff], but was not rigorously formalised. Here I seek to rectify that oversight with two kinds of contribution. Formally, I add a Stack-Time Semantics module to Stack Theory. This module includes induced layer time, window environments, temporal lifting operators, and algebraic laws relating “within-window” satisfaction to conjunction. Using this, I refine the Temporal Gap into a testable architectural question about synchrony and concurrency.
@@
Beyond Stack Theory, there are many scientific theories which emphasise some form of integration or unity across time and causal organisation. For example there is global workspace theory with its broadcast dynamics [undef, undefo], integrated information [undefz, undefu], higher-order representation [undefv], attention-schema modelling [undefq], predictive-processing and active-inference formalisms [undefp], and more general analyses of time and consciousness [undefr, undefaa]. The aim of this paper is not to adjudicate among theories, but to extract a common formal issue they face. Assume for the sake of argument that phenomenal content is modelled as a conjunction of grounded ingredients. If so, what does it then mean for that conjunction to be realised within the temporal extent of a subjective moment?
@@ 1.1 Results
    Compositional grounding. I prove that any higher-layer statement can be grounded compositionally down to the base layer while preserving truth conditions (Theorem 1).
@@
    Subjective vs objective time. I formalise higher-layer ticks as maximal objective-time blocks of constant encoding, and show they can be arbitrarily sparse (Proposition 1).
@@
    Temporal lifting algebra. I define finite-window temporal lifting operators and prove the core non-commutation result underpinning the Temporal Gap (Theorem 3).
@@
    StrongSync vs WeakSync postulates. I state both postulates in terms of a higher-layer moment statement and its base-layer grounding (Definitions 19 and 20).
@@
    Concurrency capacity threshold. I show that architectures with limited simultaneous contribution can satisfy ingredient-wise occurrence while forbidding objective co-instantiation beyond capacity (Theorem 5).
@@
    A short argument supporting StrongSync. I briefly connect StrongSync to evidence linking consciousness to synchrony and effective connectivity (Section 7).
@@ 2 Stack Theory Primitives
I follow the core notation of Stack Theory [undefd, undeff].
@@
An environment is a nonempty set $\Phi$ of mutually exclusive states. A program is a set $p\subseteq\Phi$. A program $p$ is true at $\phi\in\Phi$ iff $\phi\in p$. Let $P:=2^{\Phi}$ be the set of all programs. A vocabulary is any finite $\mathfrak{v}\subseteq P$ (finite unless explicitly stated otherwise).
@@
An environment is the set of possible states. A program is a set of states, each state representing a point in time and a smallest possible change in an environment, equating time with change [undeff]. A program is true at a state (because a state is an objective time) when the state is in that set. Conjunction is the intersection of the sets.
@@
A statement is a finite bundle of programs that must all be true at once. A completion adds more programs and makes the statement more specific. The extension is the collection of all completions that still include the original statement.
@@
A task says which input statements are allowed and which output statements count as correct. A policy is correct when it allows exactly the correct outputs for the allowed inputs.
@@ 3 Grounding Stack Theory
The abstractor takes completions of the policy and turns each completion into one set of states. The collection of these sets forms the vocabulary of the next layer. It’s basically a compression algorithm that takes completions and returns a set of equivalent programs.
@@
A stack of depth $m$ is a sequence of uninstantiated task functions $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{m}\rangle$ such that for each $i<m$, $\lambda^{i+1}$ is a restriction of $\lambda^{i}$ (written $\lambda^{i+1}\sqsubset\lambda^{i}$). Formally, each $\lambda^{i}$ assigns to each vocabulary $\mathfrak{v}$ a task $\lambda^{i}(\mathfrak{v})$ as in Definition 3. The notation $\Pi_{\lambda^{i}(\mathfrak{v})}$ denotes the set of policies correct for that task. A stack state for such a stack is a pair of sequences
@@
At each layer you pick a policy that solves the task for that layer. That policy also defines the next vocabulary, which is the higher level way of describing what is going on.
@@ 3.1 Grounding a higher-layer into lower layers
Let $\mathfrak{v}\subseteq P$ and $\pi\in\mathcal{L}_{\mathfrak{v}}$, and set $\mathfrak{v}^{\prime}=\mathfrak{f}(\mathfrak{v},\pi)$. Then for every program $p\in\mathfrak{v}^{\prime}$ there exists a completion $o_{p}\in\mathrm{Ext}(\pi)$ such that $\mathrm{T}(o_{p})=p$. Moreover, for any statement $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$ there exists a statement $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$ with
$\mathrm{T}(l)=\mathrm{T}(\mathrm{ground}_{\pi}(l)).$
Proof.
By Definition 4, for each $p\in\mathfrak{v}^{\prime}$ there exists $o_{p}\in\mathrm{Ext}(\pi)$ with $\bigcap_{r\in o_{p}}r=p$, hence $\mathrm{T}(o_{p})=p$. Fix $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$. For each $p\in l$, choose such a witness completion $o_{p}$ and define
$\mathrm{ground}_{\pi}(l):=\bigcup_{p\in l}o_{p}\subseteq\mathfrak{v}.$
Then
$\mathrm{T}(\mathrm{ground}_{\pi}(l))$ $=\bigcap_{q\in\cup_{p\in l}o_{p}}q$
$=\bigcap_{p\in l}\bigcap_{q\in o_{p}}q$
$=\bigcap_{p\in l}\mathrm{T}(o_{p})$
$=\bigcap_{p\in l}p$
$=\mathrm{T}(l).$
Since $\mathrm{T}(l)\neq\emptyset$, also $\mathrm{T}(\mathrm{ground}_{\pi}(l))\neq\emptyset$. Hence $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$. ∎
Proof intuition.
Each macro-ingredient has a lower-layer completion with the same truth set. Replacing ingredients and intersecting preserves the overall truth conditions.
@@ 3.2 Compositional grounding
Grounding works like compiling a high level statement into lower level pieces. Replace each high level ingredient with a lower level statement that means the same thing. Repeating this down the stack gives a base level conjunction that is true in exactly the same cases.
@@
Proof intuition.
Iterate Lemma 1 down the stack and every grounding step preserves the truth set, so the composed grounding does too.
Remark 1 (Non-uniqueness (and why that is acceptable)).
Grounding is generally non-unique because the witness completions $o_{p}$ in Lemma 1 need not be unique. For Temporal Gap questions, what matters is that some base-layer grounded conjunction with the same truth set exists. Different groundings correspond to different decompositions of the same higher-layer content into base-level ingredients.
@@ 4 Stack-Time Semantics Module
This section is designed as a reusable addendum to Stack Theory. It defines induced layer time, window semantics, and temporal lifting operators in a way that can be cited by future work.
@@ 4.1 Objective time and layer time as a quotient
Objective time is represented by a discrete index $t\in\mathbb{N}$. An objective trajectory is a function $\tau:\mathbb{N}\to\Phi$ such that
$\tau(t{+}1)\neq\tau(t)\quad\text{for all }t\in\mathbb{N}.$
@@
Objective time is a step counter for objective change. If nothing changes then there is no next objective index. Two events (statements) are simultaneous when they occur at the same objective index.
@@
For a vocabulary $\mathfrak{v}\subseteq P$ and a state $\phi\in\Phi$, define
$\mathrm{Enc}_{\mathfrak{v}}(\phi):=\{\,p\in\mathfrak{v}:\phi\in p\,\}\subseteq\mathfrak{v}.$
@@
Given a state and a vocabulary, the encoding collects all programs in that vocabulary that are true at that state. This gives the most detailed statement you can write in that vocabulary that still holds at that state.
@@
Proof.
Since $\phi\in p$ for every $p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)$, I have $\phi\in\bigcap_{p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)}p=\mathrm{T}(\mathrm{Enc}_{\mathfrak{v}}(\phi))$, so the intersection is nonempty and $\mathrm{Enc}_{\mathfrak{v}}(\phi)\in\mathcal{L}_{\mathfrak{v}}$. If $l$ is true at $\phi$, then $\phi\in p$ for all $p\in l$, hence $p\in\mathrm{Enc}_{\mathfrak{v}}(\phi)$ for all $p\in l$, so $l\subseteq\mathrm{Enc}_{\mathfrak{v}}(\phi)$. ∎
Proof intuition.
Every program in $\mathrm{Enc}_{\mathfrak{v}}(\phi)$ contains $\phi$, so their intersection contains $\phi$ and any statement true at $\phi$ can only use such programs.
@@
Given an objective trajectory $\tau$ and a vocabulary $\mathfrak{v}$, define the induced layer trajectory
$\tau^{(\mathfrak{v})}(t):=\mathrm{Enc}_{\mathfrak{v}}(\tau(t))\in\mathcal{L}_{\mathfrak{v}}.$
Define the layer-tick map $\kappa_{\mathfrak{v}}:\mathbb{N}\to\mathbb{N}$ by $\kappa_{\mathfrak{v}}(0)=0$ and, for $t>0$,
@@
Define an equivalence relation on objective times by $t\approx_{\mathfrak{v}}t^{\prime}$ iff $\kappa_{\mathfrak{v}}(t)=\kappa_{\mathfrak{v}}(t^{\prime})$. The corresponding layer time is the quotient set $\mathbb{N}/\!\approx_{\mathfrak{v}}$. Each equivalence class is a maximal contiguous block of objective times on which the induced macrostatement $\tau^{(\mathfrak{v})}$ is constant.
@@
For each objective time step, write down the encoding at that step. If the encoding stays the same for several steps, treat those steps as one higher level moment. The tick counter increases only when the encoding changes, so higher level time advances only when something visible at that level changes.
@@
For any $N\in\mathbb{N}$ there exist $\Phi,\mathfrak{v},\tau$ such that the induced trajectory $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps. The objective-time gap between successive layer changes can be arbitrarily large.
Proof.
Let $\Phi=\{\phi_{t}:t\in\mathbb{N}\}$. Let $p:=\{\phi_{0},\phi_{1},\dots,\phi_{N}\}$ and take $\mathfrak{v}=\{p\}$. Define $\tau(t)=\phi_{t}$ for all $t\in\mathbb{N}$. Then $\tau$ is an objective trajectory because $\tau(t{+}1)\neq\tau(t)$ for all $t$. For all $t\leq N$ I have $\tau(t)\in p$, so $\tau^{(\mathfrak{v})}(t)=\{p\}$. I also have $\tau(N{+}1)\notin p$, so $\tau^{(\mathfrak{v})}(N{+}1)=\emptyset$. Hence $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps. ∎
Proof intuition.
Choose a vocabulary that cannot distinguish many distinct objective microstates. The induced macrostate then stays constant across many objective changes, so layer time barely advances.
@@ 4.2 Time and lifting operators
A subjective moment is commonly modelled as having nonzero temporal extent (the “specious present”) [undefr, undefaa]. To speak precisely about satisfaction within such a window, I introduce windowing maps and their induced temporal operators.
Definition 10 (Windowing maps).
A windowing map is a function $\mathrm{W}:\mathbb{N}\to\mathcal{P}(\mathbb{N})$ such that $\mathrm{W}(t)$ is finite and nonempty for every $t$. For horizon $\Delta\in\mathbb{N}$ and stride $s\in\mathbb{N}_{>0}$, the sliding window map is
$\mathrm{W}_{\Delta,s}(t)=\{st,st{+}1,\dots,st{+}\Delta\}.$
Intuition.
A windowing map picks, for each subjective index, a small set of objective time steps that belong to that moment. The parameter $\Delta$ controls how far the window extends, so the window contains $\Delta{+}1$ time steps. The parameter $s$ controls how far you shift the window each time.
@@
Fix a windowing map $\mathrm{W}$, an objective trajectory $\tau$, and a subjective index $t$. For a program $p\subseteq\Phi$, define
@@
Ingredient-wise means every ingredient becomes true at least once somewhere in the window. Co-instantiated means there is some time step in the window where all ingredients are true together.
@@
For conjunctive $l$, ingredient-wise satisfaction is of the form $\forall p\in l\ \exists u\in\mathrm{W}(t)$, while co-instantiation is $\exists u\in\mathrm{W}(t)\ \forall p\in l$. The Temporal Gap is precisely the possible separation between these quantifier orders.
@@
Fix $\Delta\in\mathbb{N}$. Define the window environment $\Phi^{\Delta}:=\Phi^{\{0,1,\dots,\Delta\}}\cong\Phi^{\Delta+1}$ whose elements are windows $\sigma=(\phi_{0},\dots,\phi_{\Delta})$. Given stride $s$ and an objective trajectory $\tau$, define the induced window trajectory
$\tau^{\Delta,s}(t):=\big(\tau(st),\tau(st{+}1),\dots,\tau(st{+}\Delta)\big)\in\Phi^{\Delta}.$
@@
Treat a whole window of $\Delta{+}1$ consecutive states as a single new state. This lets you test lifted statements one window at a time.
@@
$\Diamond_{\Delta}p$ means $p$ is true at least once somewhere in the window. $\Box_{\Delta}p$ means $p$ is true at every time step in the window.
@@
To lift a statement, lift each ingredient and then require all lifted ingredients to hold for the window. With $\Diamond_{\Delta}$ this means every ingredient shows up somewhere in the window. With $\Box_{\Delta}$ this means every ingredient stays true throughout the window.
Remark 3 (Internalising temporal semantics as an abstraction layer).
The window environment $\Phi^{\Delta}$ and the lifted vocabulary $\{\Diamond_{\Delta}p:p\in\mathfrak{v}\}$ (or $\{\Box_{\Delta}p\}$) together define a new abstraction layer in the native Stack-Theory sense. This is the formal bridge between “over time” constraints and pointwise Stack-Theory semantics.

### v2

@@ Abstract
Whether machines can be conscious depends not only on what they compute, but when they compute it. Most deployed artificial systems realise their functions via sequential or time-multiplexed updates, yet a moment of conscious experience feels unified and simultaneous. I prove that this difference matters. I augment Stack Theory with algebraic laws relating within time-window constraint satisfaction to conjunction. I introduce a temporal semantics over windowed trajectories $\tau_{\Delta}$ and prove that existential temporal realisation $\Diamond_{\Delta}$ does not preserve conjunction. A system can realise all the ingredients of experience across time without ever instantiating the experienced conjunction itself. I then distinguish two postulates, Chord and Arpeggio. Chord is the position that conscious unity requires objective co-instantiation of the grounded conjunction within the window, like a musical chord. Arpeggio only needs the ingredients to occur within window, like a melody. I formalise concurrency-capacity to measure what is needed to satisfy co-instantiation. Finally, I review neurophysiological evidence suggesting that consciousness depends on phase synchrony and effective connectivity, and that loss of consciousness is associated with its breakdown. Under Chord, software consciousness on strictly sequential substrates is impossible for contents whose grounding requires two or more simultaneous contributors. The hardware matters.
@@ 1 Introduction
A central open question in machine consciousness is whether behaviourally and functionally matched systems must be matched in phenomenal realisation. Stack Theory approaches this by formalising cognitive processes in terms of abstraction layers [7, 4, 2]. Various aspects of incumbent theories of consciousness are formalised [24, 12, 16, 23, 26, 1], organised into a hierarchy of causal identities [4], and derived as a necessary consequence of a generalisation-optimal learning theory that makes valence primary [6, 8, 5, 3]. Those necessary ingredients of consciousness are beyond the scope of this paper. Here I am concerned with a question of sufficiency. Stack Theory shows how the behaviour of two systems can appear identical at one level of abstraction, while being different in the levels below. Conscious experience can be purely functional, but still depend on more than surface level function—not just on what is computed, but when and how. The ingredients of what looks like a single experienced moment at a high level of abstraction may be smeared across objective micro-time at a lower level, such that no objective time-slice contains the full grounded conjunction. If the ingredients of a subjective moment need not synchronise in objective time, then current hardware suffices to host machine consciousness—and a “liquid-brained” swarm intelligence [25] like an ant colony can be conscious. But if synchronised broadcast and causal co-dependence matter, then far fewer systems qualify. This question of architectural sufficiency is the Temporal Gap.
The Temporal Gap was proposed in earlier work [7], but was not rigorously formalised. Here I add a Stack-Time Semantics module to Stack Theory, including induced layer time, window environments, temporal lifting operators, and algebraic laws relating “within-window” satisfaction to conjunction. Using this, I refine the Temporal Gap into a testable architectural question about synchrony and concurrency.
@@
Beyond Stack Theory, many scientific theories emphasise some form of integration or unity across time and causal organisation: global workspace theory [1, 17], integrated information [27, 22], higher-order representation [23], attention-schema modelling [29], predictive-processing and active-inference [18], and more general analyses of time and consciousness [19, 28]. The aim here is not to adjudicate among theories, but to extract a common formal issue: if phenomenal content is modelled as a conjunction of grounded ingredients, what does it mean for that conjunction to be realised within the temporal extent of a subjective moment?
@@ 1.1 Results
    Compositional grounding. Any higher-layer statement can be grounded compositionally to the base layer preserving truth conditions (Theorem 1).
@@
    Subjective vs objective time. Higher-layer ticks are maximal objective-time blocks of constant encoding, and can be arbitrarily sparse (Proposition 1).
@@
    Temporal lifting algebra. The core non-commutation result underpinning the Temporal Gap (Theorem 3).
@@
    Chord vs Arpeggio postulates. Both postulates stated in terms of a higher-layer moment statement and its base-layer grounding (Definitions 16 and 17).
@@
    Concurrency capacity threshold. Architectures with limited simultaneous contribution can satisfy ingredient-wise occurrence while forbidding co-instantiation beyond capacity (Theorem 4).
@@
    Evidence supporting Chord. Co-instantiation requirements are connected to evidence linking consciousness to synchrony and effective connectivity (Section 7).
@@ 2 Stack Theory Primitives
I follow the core notation of Stack Theory [9, 7].
@@
An environment is a nonempty set $\Phi$ of mutually exclusive states. A program is a set $p\subseteq\Phi$. A program $p$ is true at $\phi\in\Phi$ iff $\phi\in p$. Let $P:=2^{\Phi}$ be the set of all programs. A vocabulary is any finite $\mathfrak{v}\subseteq P$.
@@
An environment is the set of possible states. A program is a set of states, each state representing a point in time and a smallest possible change in an environment, equating time with change [7]. A program is true at a state when the state is in that set. Conjunction is set intersection.
@@
A statement is a finite bundle of programs that must all be true at once. A completion adds more programs, making the statement more specific. The extension collects all completions that still include the original statement.
@@
A task specifies which inputs are allowed and which outputs count as correct. A policy is correct when it selects exactly the correct outputs for the allowed inputs.
@@ 3 Grounding Stack Theory
The abstractor takes completions of the policy and turns each into one set of states. The collection of these sets forms the vocabulary of the next layer—a compression that maps completions to equivalent programs.
@@
A stack of depth $m$ is a sequence of uninstantiated task functions $\langle\lambda^{0},\lambda^{1},\dots,\lambda^{m}\rangle$ such that for each $i<m$, $\lambda^{i+1}$ is a restriction of $\lambda^{i}$ (written $\lambda^{i+1}\sqsubset\lambda^{i}$). Each $\lambda^{i}$ assigns to each vocabulary $\mathfrak{v}$ a task $\lambda^{i}(\mathfrak{v})$. The notation $\Pi_{\lambda^{i}(\mathfrak{v})}$ denotes the set of correct policies for that task. A stack state is a pair of sequences
@@
At each layer you pick a policy that solves the task for that layer. That policy also defines the next vocabulary, the higher level description of what is going on.
@@ 3.1 Grounding a Higher-Layer into Lower Layers
Let $\mathfrak{v}\subseteq P$ and $\pi\in\mathcal{L}_{\mathfrak{v}}$, and set $\mathfrak{v}^{\prime}=\mathfrak{f}(\mathfrak{v},\pi)$. Then for every program $p\in\mathfrak{v}^{\prime}$ there exists a completion $o_{p}\in\mathrm{Ext}(\pi)$ such that $\mathrm{T}(o_{p})=p$. Moreover, for any statement $l\in\mathcal{L}_{\mathfrak{v}^{\prime}}$ there exists a statement $\mathrm{ground}_{\pi}(l)\in\mathcal{L}_{\mathfrak{v}}$ with $\mathrm{T}(l)=\mathrm{T}(\mathrm{ground}_{\pi}(l))$.
Proof. See Supplementary Material. Each macro-ingredient has a lower-layer completion with the same truth set; replacing ingredients and intersecting preserves truth conditions. ∎
@@ 3.2 Compositional Grounding
Grounding works like compiling a high level statement into lower level pieces. Replace each high level ingredient with a lower level statement that means the same thing. Repeating this down the stack gives a base level conjunction true in exactly the same cases.
@@ 4 Stack-Time Semantics Module
This section defines induced layer time, window semantics, and temporal lifting operators as a reusable addendum to Stack Theory.
@@ 4.1 Objective Time and Layer Time as a Quotient
Objective time is represented by a discrete index $t\in\mathbb{N}$. An objective trajectory is a function $\tau:\mathbb{N}\to\Phi$ such that $\tau(t{+}1)\neq\tau(t)$ for all $t\in\mathbb{N}$.
@@
Objective time is a step counter for objective change. If nothing changes then there is no next objective index. Two events are simultaneous when they occur at the same objective index.
@@
For a vocabulary $\mathfrak{v}\subseteq P$ and a state $\phi\in\Phi$, define $\mathrm{Enc}_{\mathfrak{v}}(\phi):=\{\,p\in\mathfrak{v}:\phi\in p\,\}\subseteq\mathfrak{v}$.
@@
Given a state and a vocabulary, the encoding collects all programs in that vocabulary true at that state—the most detailed true statement in that vocabulary.
@@
Proof. See Supplementary Material. ∎
@@
Given an objective trajectory $\tau$ and a vocabulary $\mathfrak{v}$, define the induced layer trajectory $\tau^{(\mathfrak{v})}(t):=\mathrm{Enc}_{\mathfrak{v}}(\tau(t))\in\mathcal{L}_{\mathfrak{v}}$. Define the layer-tick map $\kappa_{\mathfrak{v}}:\mathbb{N}\to\mathbb{N}$ by $\kappa_{\mathfrak{v}}(0)=0$ and, for $t>0$,
@@
Define $t\approx_{\mathfrak{v}}t^{\prime}$ iff $\kappa_{\mathfrak{v}}(t)=\kappa_{\mathfrak{v}}(t^{\prime})$. The layer time is $\mathbb{N}/\!\approx_{\mathfrak{v}}$: each equivalence class is a maximal contiguous block of objective times on which $\tau^{(\mathfrak{v})}$ is constant.
@@
For each objective time step, write down the encoding. If it stays the same for several steps, treat them as one higher level moment. Layer time advances only when something visible at that level changes.
@@
For any $N\in\mathbb{N}$ there exist $\Phi,\mathfrak{v},\tau$ such that the induced trajectory $\tau^{(\mathfrak{v})}$ changes at most once in the first $N$ objective steps.
Proof. See Supplementary Material. Choose a vocabulary that cannot distinguish many distinct objective microstates; the macrostate stays constant. ∎
@@ 4.2 Time and Lifting Operators
A subjective moment is commonly modelled as having nonzero temporal extent (the “specious present”) [19, 28]. I introduce temporal operators to formalise satisfaction within such a window.
@@
Fix a windowing map $\mathrm{W}$, an objective trajectory $\tau$, and a window index $t$. For a program $p\subseteq\Phi$, define
@@
Ingredient-wise means every ingredient becomes true at least once somewhere in the window. Co-instantiated means there is a time step in the window where all ingredients are true together.
@@
For conjunctive $l$, ingredient-wise satisfaction is $\forall p\in l\ \exists u\in\mathrm{W}(t)$, while co-instantiation is $\exists u\in\mathrm{W}(t)\ \forall p\in l$. The Temporal Gap amounts to whether conscious unity requires ingredient-wise satisfaction or co-instantiation.
@@
Fix $\Delta\in\mathbb{N}$. Define the window environment $\Phi^{\Delta}:=\Phi^{\{0,1,\dots,\Delta\}}\cong\Phi^{\Delta+1}$ whose elements are windows $\sigma=(\phi_{0},\dots,\phi_{\Delta})$. Given an objective trajectory $\tau$, define the induced window trajectory $\tau_{\Delta}(t):=\big(\tau(t),\tau(t{+}1),\dots,\tau(t{+}\Delta)\big)\in\Phi^{\Delta}$.
@@
Treat a whole window of $\Delta{+}1$ consecutive states as a single new state.
@@
$\Diamond_{\Delta}p$ means $p$ is true at least once in the window. $\Box_{\Delta}p$ means $p$ is true at every step in the window.
@@
To lift a statement, lift each ingredient and require all lifted ingredients to hold. With $\Diamond_{\Delta}$ this means every ingredient shows up somewhere in the window.
Remark 2 (Temporal semantics as an abstraction layer).
The window environment $\Phi^{\Delta}$ and the lifted vocabulary $\{\Diamond_{\Delta}p:p\in\mathfrak{v}\}$ define a new abstraction layer in the native Stack-Theory sense, bridging “over time” constraints and pointwise semantics. The compilation of history-dependent (pathwise) abstractors into pointwise abstractors on augmented state is detailed in the Supplementary Material.